Two cars travel the same distance. The first car travels at a rate of 40 miles per hour and reaches its destination in t hours. The second car travels at a rate of 55 miles per hour and reaches its destination 3 hours earlier than the first car. How long does it take for the first car to reach its destination? How long does it take for the second car to reach its destination?

Answer by mananth(16946)

Two cars travel the same distance. The first car travels at a rate of 40 miles per hour and reaches its destination in t hours. The second car travels at a rate of 55 miles per hour and reaches its destination 3 hours earlier than the first car. How long does it take for the first car to reach its destination? How long does it take for the second car to reach its destination?
...
Two cars travel the same distance.
The first car rate of 40 miles per hour t hours.
The second car rate of 55 miles per hour 3 hours earlier than the first car.
distance = rate * time
First Car
D= 40t
Second car D=55*(t-3)
..
40t=55(t-3)
40t=55t-165
-55t
40t-55t=-165
-15t=-165
/-15
t=11 hours First car
Second Car takes 11-3=8 hours
...
m.ananth@hotmail.ca